a certain gas expands in volume from 2.0 L to 6.0 L at constant temperature.


calculate the work done by the gas if it expands:


a) against a vacuum


b) against a constant pressure of 1.2 atm


1
Expert's answer
2021-04-21T07:23:33-0400

a)W = P∆V

= 6 - 2 = 4

= 4 Joules

b) W = P∆V

= 1.2(6-2)

= 4 × 1.2

= 4.8Joules